OpenStudy (anonymous):

Let w=f(x,y) be a function of two independent variables x,y. Write sown a vector which is perpendicular to the tangent plane at the point (x0,y0,z0). What formula could I use here? Or how would I go about solving this, I know that the result would be that the vector which is normal to the surface and so the normal vector dotted with any vector in the plane is 0

OpenStudy (experimentx):

looks like gradient problem.

OpenStudy (experimentx):

let's try for a sphere ... problem will be greatly simplified.

OpenStudy (anonymous):

ooh er sphere is an implicit isnt it?

OpenStudy (anonymous):

how about a parabaloid: z = x^2 + y^2

OpenStudy (experimentx):

okay that should do fine ... wow ... it's called parabaloid!!

OpenStudy (anonymous):

I have the equation for the tangent plane; \[z=f(x _{0},y _{0}) +[df/dx (x _{0},y _{0})] (x-x_{0})+[df/dy (x _{0},y _{0})] (y-y_{0})\] and I read somewhere that the normal vector to a tangent plane is the gradient of the tangent plane which would give; \[([df/dx(x_{0},y_{0})],[df/dy(x_{0},y_{0})],-1)\] but that doesn't make sense

OpenStudy (anonymous):

i know right! i found out the other day and i cant stop saying it

OpenStudy (anonymous):

parabaloid

OpenStudy (experimentx):

V = 2x i + 2 y j +k is this going to be normal to z = x^2 + y^2 or not ..??

OpenStudy (anonymous):

I honestly don't know :(

OpenStudy (experimentx):

anyone know matlab??

OpenStudy (anonymous):

answer mark 3 let us have function f(x,y) let partial derivatives be \[f'_x(x) \text{ , } f'_y(y)\] so we can find tangents parallel to x axis and y axis eg. x tangent vector, \[v_x\] will be in the y=0 plane and will have gradient of\[ f'_x(x_0)\] so will be parrallel to \[\left(\begin{matrix}1 \\ 0 \\f'_x(x_0)\end{matrix}\right)\] likewise \[v_y = \left(\begin{matrix}0 \\ 1 \\f'_y(y_0)\end{matrix}\right)\] so computing their cross product we get \[\pm\left(\begin{matrix}f'_x(x_0) \\ f'_y(y_0)\\1\end{matrix}\right)\] as a normal (plus or minus for the order you cross them)

OpenStudy (anonymous):

i know of matlab, i dont have it. is it good?

OpenStudy (anonymous):

I know matab

OpenStudy (anonymous):

matlab*

OpenStudy (experimentx):

know how to plot 3d vector lines at points??

OpenStudy (anonymous):

for z = x^2 + y^2 we have \[2x_0i + 2y_0j - k\] i think

OpenStudy (anonymous):

as with all my answers on this these are just educated guesses at best

OpenStudy (anonymous):

Thats what I had @eigenschmeigen and the way you've explained it makes sense, @experimentX I don't know how to do that, I've only done one course in it which was mostly project based

OpenStudy (experimentx):

yeah ... @eigenschmeigen thats amaizing.

OpenStudy (turingtest):

yeah, @eigenschmeigen has the right answer, and it's not too hard as experiment said in the beginning it's a gradient problem

OpenStudy (experimentx):

my vector should have been V = 2x i + 2 y j - k I realize now ... lol

OpenStudy (experimentx):

@TuringTest it was quite a worthy to prove that gradient will be perpendicular to surface.

OpenStudy (turingtest):

just define\[f(x,y,z)=x^2+y^2-z\]\[\nabla f=<2x,2y,-1>\]and evaluate it at whatever point... and yeah @experimentX that's a nice thing to be able to prove

OpenStudy (anonymous):

whats the name of the delta symbol here? does it have a name?

OpenStudy (turingtest):

"del" it is usually called

OpenStudy (anonymous):

is that delta?

OpenStudy (turingtest):

"del of f"

OpenStudy (experimentx):

actually i was looking for it myself. It's called Del operator and it's equitvalent to (id/dx+jd/dy+kd/dz)

OpenStudy (anonymous):

oh cool

OpenStudy (turingtest):

\[\nabla f\]\[\nabla\times f\]\[\nabla\cdot f\]gradient, curl, and divergence respectively

OpenStudy (turingtest):

though in latex it's called a "nabla" I don't know why...

OpenStudy (anonymous):

wait its called gradient but it gives you the normal vector? weird

OpenStudy (experimentx):

Yup ... there's an interesting theorem on this.

OpenStudy (turingtest):

\[\nabla f\] is the fastest possible rate change direction of f if f describes a surface, like in this problem, it should make sense that the fastest possible rate change of f is directly out of the shape, i.e. perpendicular to it if we instead took\[\nabla z\](imagine it's the equation f a hill or something) we would get a vector that points along the parabaloid in the "steepest" direction possible

OpenStudy (anonymous):

ahh i see

OpenStudy (anonymous):

GRAONFKAMKASMDKLASMl just found a mistype in my answer

OpenStudy (turingtest):

I didn't o-0

OpenStudy (anonymous):

the cross product, the k component should be negative..

OpenStudy (anonymous):

its really really annoying me

OpenStudy (anonymous):

z component

OpenStudy (turingtest):

oh yeah, I see now signs get lost easily in cross-products :P it happens...

OpenStudy (anonymous):

\[\left(\begin{matrix}f'_x(x_0) \\ f'_y(y_0)\\-1\end{matrix}\right)\]